# Triplet Interaction Improves Graph Transformers: Accurate Molecular Graph Learning with Triplet Graph Transformers

### v1

@@ Abstract
Graph transformers typically lack direct pair-to-pair communication, instead forcing neighboring pairs to exchange information via a common node. We propose the Triplet Graph Transformer (TGT) that enables direct communication between two neighboring pairs in a graph via novel triplet attention and aggregation mechanisms. TGT is applied to molecular property prediction by first predicting interatomic distances from 2D graphs and then using these distances for downstream tasks. A novel three-stage training procedure and stochastic inference further improve training efficiency and model performance. Our model achieves new state-of-the-art (SOTA) results on open challenge benchmarks PCQM4Mv2 and OC20 IS2RE. We also obtain SOTA results on QM9, MOLPCBA, and LIT-PCBA molecular property prediction benchmarks via transfer learning. We also demonstrate the generality of TGT with SOTA results on the traveling salesman problem (TSP). Our code is available at https://github.com/shamim-hussain/tgt.
@@ 1 Introduction
Graph transformers utilize global self-attention to enable dynamic information exchange among node representations. Additionally, since graph topology and edge representations are as crucial as node representations for many tasks, the Edge-augmented Graph Transformer (EGT) (Hussain et al., 2022) introduces dedicated edge channels that are updated across layers, enabling new pairwise (i.e., both existing and non-existing edge) representations to emerge over consecutive layers. This explicit modeling of both node and edge embeddings benefits performance on both node-centric and pairwise/link prediction tasks. However, although EGT allows information flow between node and pair representations, it lacks direct communication between pairs. Instead, neighboring pairs can only exchange information via their common node, creating a bottleneck. Direct edge-to-edge communication is critical where the representation of an edge is strongly correlated with edges and . This is particularly important for geometric graphs where constraints like triangle inequality must be followed.
@@
In particular, 3D molecular geometry plays a vital role in determining chemical properties. While molecular graphs represent atoms as nodes and bonds as edges (“2D” structure), the relative positions of atoms in 3D space crucially influence quantum mechanical attributes like orbital energies and dipole moments and also other properties like solubility and interactions with proteins. Accordingly, prior works (Stärk et al., 2022; Liu et al., 2021b) have shown that incorporating 3D structure improves performance on molecular property prediction. However, determining ground truth 3D geometries requires expensive quantum chemical simulations, presenting computational barriers for large-scale inference. In this work, we explicitly learn to predict the molecular structure using only 2D topological information. Specifically, we train a model to predict interatomic distances, serving as structural input to the downstream chemical property prediction task.
We make several key contributions in terms of architecture and training for molecular property prediction and also for graph learning in general. To address the edge-node-edge bottleneck, we introduce a novel triplet interaction mechanism that enables direct communication between neighboring pairs without involving a common node. This also improves the downstream property prediction task, as it makes the model robust to structural perturbations. We call this architecture Triplet Graph Transformer (TGT). We also demonstrate the effectiveness of triplet interaction in other graph learning tasks, such as the traveling salesman problem (TSP) – demonstrating its generality.
@@
We also propose a three-stage training framework for the distance and task predictors, which significantly reduces the training time and improves the performance of the model. We also introduce a novel stochastic inference technique that further improves the performance of the model and allows for non-iterative parallel inference and uncertainty estimation. Additionally, we introduce new methods for regularizing both the pairwise update and triplet interaction mechanisms. We also propose a locally smooth structural noising method and a binned distance prediction objective that makes the model robust to structural perturbations.
Through these contributions, our proposed TGT model achieves new state-of-the-art (leaderboard) results on the PCQM4Mv2 (Hu et al., 2021), OC20 IS2RE (Chanussot et al., 2021), and QM9 (Ramakrishnan et al., 2014) quantum chemical datasets. We also demonstrate the transferability of our learned distance predictor by achieving SOTA results on the MOLPCBA (Hu et al., 2020) and LIT-PCBA (Tran-Nguyen et al., 2020) benchmarks, which are molecular property prediction and drug discovery datasets respectively. This showcases the ability of our trained distance predictor to act as an off-the-shelf pairwise feature extractor that can be utilized for new molecular graph learning tasks.
@@ 2 Related Work
Some previous works like GraphTrans (Wu et al., 2021), GSA (Wang & Deng, 2021), and GROVER (Rong et al., 2020) and some new works like GPS (Rampášek et al., 2022) and GPS++ (Masters et al., 2022) have used global self-attention to boost the expressivity of GNNs in a hybrid approach. But our work is more directly related to the recently proposed pure graph transformers (GTs) such as Graphormer (Ying et al., 2021), EGT (Hussain et al., 2022), GRPE (Park et al., 2022), GEM-2 (Liu et al., 2022a), and UniMol+ (Lu et al., 2023). Our contribution is a novel triplet interaction mechanism for direct edge-to-edge communication which improves the expressivity of GTs. We primarily approach the problem of molecular property prediction from a pure graph transformer perspective (but also demonstrate its generality). Recently, this problem has seen a lot of interest in the form of equivariant/invariant GNNs like SchNet (Schütt et al., 2017), DimeNet (Gasteiger et al., 2020b), GemNet (Gasteiger et al., 2021), SphereNet (Liu et al., 2021b) and equivariant transformers like Equiformer (Thölke & De Fabritiis, 2021), and TorchMD-Net (Thölke & De Fabritiis, 2022). Unlike these works, we preserve SE(3) invariance by limiting the input features to interatomic distances. We train our network to predict interatomic distances from 2D molecular graphs, which allows for inference even in the absence of 3D information. This is in contrast to 3D pretraining approaches like 3D Infomax (Stärk et al., 2022), GraphMVP (Liu et al., 2021a), Chemrl-GEM (Fang et al., 2021), 3D PGT (Wang et al., 2023b), GeomSSL (Liu et al., 2022b), Transformer-M (Luo et al., 2022) which resort to 2D finetuning or multitask learning to make predictions in absence of 3D information. On the other hand, UniMol+ (Lu et al., 2023) iteratively refines cheaply computed RDKit (Landrum, 2013) coordinates. In contrast, our approach does not require any initial 3D coordinates and directly predicts interatomic distances from 2D molecular graphs.
@@ 3.1 TGT Architecture
Triplet Graph Transformer (TGT) significantly extends the Edge-augmented Graph Transformer (EGT) (Hussain et al., 2022) (see Appendix C for details) by introducing direct pair-to-pair communication. While EGT maintains pairwise embeddings for every node pair , it lacks a direct mechanism for modeling interactions between two pairs. This limitation arises from updating the pairwise embeddings solely based on the node embeddings and . While this choice ensures a computational complexity of like the original transformer, it constrains the model’s expressivity. Enabling completely unconstrained pair interactions would escalate computational complexity to , which is intractable for any but the smallest graphs. A judicious compromise involves identifying a crucial bottleneck in adjacent pairs, namely those sharing a common node. Illustrated in Figure 1, this bottleneck is exemplified by the pair only being able to interact with the pair through the intermediary node . However, the pair contends with all other pairs (e.g., , ) that share the node , creating a bottleneck. This is mitigated by allowing direct interaction between pairs and without necessitating involvement of the node . The linear arrows between a pair of nodes in Figure 1 denote the information flow in the node channels, whereas the curved arrow represents the information flow between pairwise embeddings due to triplet interaction. This interaction should also consider the third arm of the triangle, i.e., the pair . Focusing on adjacent pairs sharing at least one common node limits the complexity to . Sub-cubic complexity is achievable with some concessions, as we will see in the next section. Some previous works have also addressed this problem via axial attention (Liu et al., 2022a) or triangular update (Lu et al., 2023; Jumper et al., 2021) (see Appendix B for details).
We introduce a novel triplet interaction module to the edge channels, between the pairwise attention block and the edge Feed Forward Network (FFN) block. This module follows the same pre-norm layer normalization and residual connection pattern as the rest of the network. The resultant Triplet Graph Transformer (TGT) architecture is shown in Figure 2(a). The triplet interaction module is shown in Figure 2(b). We propose two forms of interaction mechanisms called triplet attention and triplet aggregation and refer to the resultant variants as TGT-At and TGT-Ag respectively.
@@
where is derived from a learnable projection of the pairwise embedding and is the attention weight assigned to the pair by the pair . This is done for multiple attention heads and is the output of an attention head. and are the query and the key vectors derived from the pairwise embeddings and respectively, and are scalars derived from the pairwise embeddings belonging third arm of the triangle which participates by providing these bias and gating terms, respectively. The gating term is not strictly necessary but improves the performance of the model. We call this an inward update. Another parallel update, called outward update, is done by changing the direction of the aggregated pairs as follows:
Finally, and for all heads are concatenated and the the pairwise embedding is updated from a learnable projection of the resultant. Triplet attention combines the strengths of the axial attention mechanism (Liu et al., 2022a) and the triangular update (Jumper et al., 2021) in a single update. It is proven to be the most expressive form of interaction in our experiments, exceeding both of the aforementioned mechanisms. However, it has an compute complexity and an memory complexity (without a specialized GPU kernel).
@@
Notice that it is a tensor multiplication between the weight matrix and the value matrix, each of which has only elements and thus, has a subcubic complexity of (Ambainis et al., 2015) which is much better than the cubic complexity of triplet attention. As a compromise, we have to remove the dependence of the weights on the junction node . The weights are only determined by the pair , i.e., the third arm of the triangle due to removing the dot product term from the weights. Thus this process is not an attention mechanism. We also have an outward update, and the final update is done by concatenating the inward and outward updates. Note that the weights are bounded and normalized (ignoring the gating term) unlike the triangular update in UniMol+ (Lu et al., 2023). Also, the values being aggregated are vectors instead of scalars, which makes the process more expressive and efficient as only a few heads are required instead of many.
A comparison of the different triplet interaction mechanisms is shown in Table 1. Triplet attention is the most expressive but triplet aggregation is more efficient and scalable. Triplet aggregation performs better than triangular update and is also slightly more efficient, as verified in our experiments.
| Axial Att. | Tria. Update | Triplet Agg. | Triplet Att. | |
| Normalized? | ||||
| Gated? | ||||
| Attention? | ||||
| Values are | Vectors | Scalars | Vectors | Vectors |
| Weighted by | ||||
| Complexity | ||||
| Expressivity | Worst | Good | Better | Best |
@@ 3.2 Training and Inference for Molecular Graphs
Our training method for molecular property prediction consists of three stages. The first stage involves training the distance predictor which predicts interatomic distances from 2D molecular graphs. The second stage involves pretraining the task predictor which predicts molecular properties on noisy structures, and the third stage involves finetuning it on the predicted distances. The three stages are shown in Figure 3 along with the inference process.
@@
Our reasoning for predicting distances instead of coordinates is that they can be directly used by the task predictor. Defining all pairwise distances also defines all the angles and dihedrals. Distances are also invariant to translation/rotation and have a small, easily learned value range. We found the accuracy of predicting individual distances is more important than strictly conforming distances to a 3D structure. Additionally, unlike UniMol+ (Lu et al., 2023), no initial 3D structure is necessary for our approach, which is a significant advantage since the whole inference pipeline can be pure neural network-based and highly scalable. We predict binned distances instead of continuous values since distance distributions are often multimodal and skewed, and thus only fully captured by the model with a cross-entropy loss. This allows us to predict the most probable distance, i.e. the mode, which is more accurate than the mean or median predictions. Quantization noise due to binning does not affect the downstream task predictor which is robust to noise.
@@ 3.3 Stochastic Inference
| Model | #Param | Valid | Test-dev |
@@
Baseline Methods: GINE (Brossard et al., 2020), GCN (Kipf & Welling, 2016), GIN (Xu et al., 2018), DeeperGCN (Li et al., 2020), Virtual Node (–VN) (Gilmer et al., 2017), TokenGT (Kim et al., 2022), GRPE (Park et al., 2022), Graphormer (Ying et al., 2021), GraphGPS (Rampášek et al., 2022), EGT (Hussain et al., 2022), GEM-2 (Liu et al., 2022a), Transformer-M (Luo et al., 2022), GPS++ (Masters et al., 2022) and UniMol+ (Lu et al., 2023)
@@
Baseline Methods: SchNet (Schütt et al., 2017), DimeNet++ (Gasteiger et al., 2020a), SphereNet (Liu et al., 2021b), GNS (Kumar & Vantassel, 2022), Graphormer-3D (Shi et al., 2022), EquiFormer (Thölke & De Fabritiis, 2021), DRFormer (Wang et al., 2023a), Moleformer (Yuan et al., 2023),UniMol+ (Lu et al., 2023), Noisy Nodes (+NN)(Godwin et al., 2021)
| Val. | Test | |||
| Uni– | TGT– | Uni– | TGT– | |
| Split | Mol+ | At | Mol+ | At |
| ID | ||||
| MAE() | 379.5 | 381.3 | 374.5 | 379.6 |
| EwT() | 11.15 | 11.15 | 11.29 | 11.50 |
| OOD Ads. | ||||
| MAE() | 452.6 | 445.4 | 476.0 | 471.8 |
| EwT() | 6.71 | 6.87 | 6.05 | 5.70 |
| OOD Cat. | ||||
| MAE() | 401.1 | 391.7 | 398.0 | 399.0 |
| EwT() | 9.90 | 10.47 | 9.53 | 9.84 |
| OOD Both | ||||
| MAE() | 402.1 | 393.6 | 408.6 | 408.4 |
| EwT() | 6.68 | 6.80 | 6.06 | 6.17 |

### v2

@@ Abstract
Graph transformers typically lack third-order interactions, limiting their geometric understanding which is crucial for tasks like molecular geometry prediction. We propose the Triplet Graph Transformer (TGT) that enables direct communication between pairs within a 3-tuple of nodes via novel triplet attention and aggregation mechanisms. TGT is applied to molecular property prediction by first predicting interatomic distances from 2D graphs and then using these distances for downstream tasks. A novel three-stage training procedure and stochastic inference further improve training efficiency and model performance. Our model achieves new state-of-the-art (SOTA) results on open challenge benchmarks PCQM4Mv2 and OC20 IS2RE. We also obtain SOTA results on QM9, MOLPCBA, and LIT-PCBA molecular property prediction benchmarks via transfer learning. We also demonstrate the generality of TGT with SOTA results on the traveling salesman problem (TSP).
@@ 1 Introduction
Graph transformers utilize global self-attention to enable dynamic information exchange among node representations. Additionally, since graph topology and edge representations are as crucial as node representations for many tasks, the Edge-augmented Graph Transformer (EGT) (Hussain et al., 2022) introduces dedicated edge channels that are updated across layers, enabling new pairwise (i.e., both existing and non-existing edge) representations to emerge over consecutive layers. This explicit modeling of both node and edge embeddings benefits performance on both node-centric and pairwise/link prediction tasks. However, although EGT allows information flow between node and pair representations, it lacks direct communication between pairs. Instead, neighboring pairs can only exchange information via their common node, creating a bottleneck. This limits the expressivity of the model by allowing only 2nd-order interactions (Joshi et al., 2023). As shown by Li et al. (2024) 3rd-order interactions in the form of direct communication between neighboring pairs, i.e., within a 3-tuple of nodes is important for understanding important geometric concepts like angles.
@@
In particular, 3D molecular geometry plays a vital role in determining chemical properties. While molecular graphs represent atoms as nodes and bonds as edges (“2D” structure), the relative positions of atoms in 3D space crucially influence quantum mechanical attributes like orbital energies and dipole moments and also other properties like solubility and interactions with proteins. Accordingly, prior works (Stärk et al., 2022; Liu et al., 2021b) have shown that incorporating 3D geometry significantly improves performance on molecular property prediction. Geometric GNNs like GemNet (Gasteiger et al., 2021) rely on input features derived from the ground truth 3D geometry to predict molecular properties. However, determining ground truth 3D geometries requires expensive quantum chemical simulations, presenting computational barriers for large-scale inference. In this work, we explicitly learn to predict the molecular geometry using only 2D topological information. Specifically, we train a model to predict interatomic distances, serving as geometric input to the downstream chemical property prediction task. This is a paradigm shift from the traditional approach of 3D to 2D transfer learning (Stärk et al., 2022) or relying on less accurate 3D geometry (Fang et al., 2021).
We introduce a novel 3rd-order interaction mechanism in the form of triplet interactions that enable direct communication between neighboring pairwise representations. This improves the expressivity of popular 2nd-order interaction based graph transformers like EGT. Reliance on 3rd-order interactions rather than features (similar to (Li et al., 2024)) allows for (i) accurate prediction of geometry from scratch, i.e., without an initial estimate of 3D coordinates, and (ii) robustness to input geometric inaccuracies. Geometric GNNs lack these capabilities due to their direct reliance on input geometric features. We propose two forms of triplet interactions called triplet attention and triplet aggregation. We call this architecture Triplet Graph Transformer (TGT). We also demonstrate the effectiveness of triplet interaction in other geometric graph learning tasks, such as the traveling salesman problem (TSP) – demonstrating its generality.
@@
We also propose a three-stage training framework for the distance and task predictors, which significantly reduces the training time and improves the performance of the model. We also introduce a novel stochastic inference technique that further improves the model’s performance and allows for non-iterative parallel inference and uncertainty estimation. Additionally, we introduce new methods for regularizing both the pairwise update and triplet interaction mechanisms. We also propose a locally smooth structural noising method and a binned distance prediction objective that makes the model robust to structural perturbations.
Through these contributions, our proposed TGT model achieves new state-of-the-art (leaderboard) results on the PCQM4Mv2 (Hu et al., 2021), OC20 IS2RE (Chanussot et al., 2021), and QM9 (Ramakrishnan et al., 2014) quantum chemical datasets. We also demonstrate the transferability of our learned distance predictor by achieving SOTA results on the MOLPCBA (Hu et al., 2020) and LIT-PCBA (Tran-Nguyen et al., 2020) benchmarks, which are molecular property prediction and drug discovery datasets, respectively. This showcases the ability of our trained distance predictor to act as an off-the-shelf pairwise feature extractor that can be utilized for new molecular graph learning tasks.
@@ 2 Related Work
Some previous works like GraphTrans (Wu et al., 2021), GSA (Wang & Deng, 2021), and GROVER (Rong et al., 2020) and some new works like GPS (Rampášek et al., 2022) and GPS++ (Masters et al., 2022) have used global self-attention to boost the expressivity of GNNs in a hybrid approach. But our work is more directly related to the recently proposed pure graph transformers (GTs) such as SAN (Kreuzer et al., 2021), Graphormer (Ying et al., 2021), EGT (Hussain et al., 2022), GRPE (Park et al., 2022), GEM-2 (Liu et al., 2022a), and UniMol+ (Lu et al., 2023). Our contribution is introducing novel 3rd-order interaction mechanisms that improve the expressivity of GTs. We primarily approach the problem of molecular property prediction from a pure graph transformer perspective (but also demonstrate its generality). Recently, this problem has seen a lot of interest in the form of equivariant/invariant geometric GNNs like SchNet (Schütt et al., 2017), DimeNet (Gasteiger et al., 2020b), GemNet (Gasteiger et al., 2021), SphereNet (Liu et al., 2021b) and equivariant transformers like Equiformer (Thölke & De Fabritiis, 2021), and TorchMD-Net (Thölke & De Fabritiis, 2022). Unlike these works, our model can be used for both general-purpose graph representation learning and geometric deep learning (Li et al., 2024). We preserve SE(3) invariance by limiting the input features to interatomic distances. We train our network to predict interatomic distances from 2D molecular graphs, which allows for inference even in the absence of 3D information. This is in contrast to 3D pretraining approaches like 3D Infomax (Stärk et al., 2022), GraphMVP (Liu et al., 2021a), Chemrl-GEM (Fang et al., 2021), 3D PGT (Wang et al., 2023b), GeomSSL (Liu et al., 2022b), Transformer-M (Luo et al., 2022) which resort to 2D finetuning or multitask learning to make predictions in absence of 3D information. On the other hand, UniMol+ (Lu et al., 2023) iteratively refines cheaply computed RDKit (Landrum, 2013) coordinates. In contrast, our approach requires no initial 3D coordinates and directly predicts interatomic distances from 2D graphs.
@@ 3.1 TGT Architecture
Triplet Graph Transformer (TGT) significantly extends the Edge-augmented Graph Transformer (EGT) (Hussain et al., 2022) by introducing direct pair-to-pair communication in the form of triplet (3rd-order) interactions. In each layer, EGT maintains both node embedding for each of the nodes and pairwise embedding for each of the node pairs (see Appendix C for details). Triplet interaction operates only on the pairwise embeddings .
TGT addresses the following important limitation of EGT – it updates the pairwise embeddings solely based on the node embeddings and . While this choice ensures a computational complexity of like the original transformer, it constrains the model’s expressivity to that of a 1-GWL test (Joshi et al., 2023). To improve upon this, we must move beyond 2nd-order interactions within the 2-tuple of nodes and consider 3rd-order interactions involving the 3-tuple of nodes . Triplet interaction allows direct information flow to the pair from the neighboring pair . To complete the 3-tuple, it also considers the pair . As illustrated in Figure 1, the linear arrows between a pair of nodes denote the information flow in the node channels, whereas the curved arrow represents the information flow between pairwise embeddings due to triplet interaction. This interaction allows to aggregate all neighboring pairs without involving the junction node , resolving the bottleneck at . This 3rd-order interaction elevates the model’s expressivity beyond 1-GWL, approaching that of 2-GWL, allowing it to model complex geometric relationships in graphs. But this comes at the cost of increased computational complexity. We only consider 3rd-order interactions which limits the complexity to . This is a good engineering choice since, as shown by Li et al. (2024), 3rd-order interactions are crucial for geometric understanding, yet 4th-order interactions add little/no benefit at much higher computational cost. Also, sub-cubic complexity is achievable with some concessions, as we will see in the next section. Some previous works have also used 3rd-order interactions via axial attention (Liu et al., 2022a) or triangular update (Lu et al., 2023; Jumper et al., 2021) (see Appendix B for details).
We introduce a novel triplet interaction module to the edge channels, between the pairwise attention block and the edge Feed Forward Network (FFN) block. This module follows the same pre-norm layer normalization and residual connection pattern as the rest of the network. The resultant Triplet Graph Transformer (TGT) architecture is shown in Figure 2(a). The triplet interaction module is shown in Figure 2(b). We propose two forms of interaction mechanisms called triplet attention and triplet aggregation and refer to the resultant variants as TGT-At and TGT-Ag, respectively.
@@
where the value vector is derived from a learnable projection of the pairwise embedding and is the attention weight assigned to the pair by the pair . This is done for multiple attention heads and is the output of an attention head. and are the query and the key vectors derived from the pairwise embeddings and , respectively. are scalars derived from the pairwise embeddings belonging to the third pair within the 3-tuple , which participates by providing these bias and gating terms, respectively. The gating term is not strictly necessary but improves the performance of the model. We call this an inward update. Another parallel update, called outward update, is done by changing the direction of the aggregated pairs as follows:
Finally, and for all heads are concatenated and the the pairwise embedding is updated from a learnable projection of the resultant. Triplet attention combines the strengths of axial attention (Liu et al., 2022a) and triangular update (Jumper et al., 2021) in a single update. It is thus the most expressive form of interaction and outperforms both of the aforementioned mechanisms in our experiments. However, it has an compute complexity.
@@
Notice that it is a tensor multiplication between the weight matrix and the value matrix, each of which has only elements and thus, has a subcubic complexity of (Ambainis et al., 2015) which is much better than the cubic complexity of triplet attention. As a compromise, we have to remove the dependence of the weights on the junction node . The weights are only determined by the pair , due to removing the dot product term from the weights. Thus, this process is not an attention mechanism. We also have an outward update, and the final update is done by concatenating the inward and outward updates. Note that the weights are bounded and normalized (ignoring the gating term) unlike the triangular update in UniMol+ (Lu et al., 2023). Also, the values being aggregated are vectors instead of scalars, which makes the process more expressive and efficient as only a few heads are required instead of many.
A comparison of different 3rd-order interaction mechanisms (previous methods and our triplet interactions) is shown in Table 1. Triplet attention is the most expressive because all other interactions are ablated versions of it, i.e., they can be derived by removing some terms from triplet attention (see Equations 1, 2, 3, 4, 5, 6, 8, 9 and 10). Triplet aggregation is more expressive than triangular update because it aggregates vector values rather than scalar values. In our experiments, it outperforms triangular update while being more efficient. Axial attention is the least expressive because it does not include the pair , and thus interaction within the 3-tuple is incomplete. In summary, triplet attention is the best-performing yet heavyweight method while triplet aggregation is more efficient and scalable at the cost of some performance. Both are better than previous methods.
@@ 3.2 Training and Inference for Molecular Graphs
Our training method for molecular property prediction consists of three stages. The first stage involves training the distance predictor which predicts interatomic distances from 2D molecular graphs. The second stage involves pretraining the task predictor which predicts molecular properties on noisy 3D structures, and the third stage involves finetuning it on the predicted distances. The three stages are shown in Figure 3 along with the inference process.
@@
Our reasoning for predicting distances instead of coordinates is that they can be directly used by the task predictor. We predict and utilize the full distance matrix which fully defines the geometry of the molecule up to a reflection (i.e., it cannot capture chirality). Thus, defining all pairwise distances also defines all the angles and dihedrals. Distances are also invariant to translation/rotation and have a small, easily learned value range. We found the accuracy of predicting individual distances is more important than strictly conforming distances to a 3D structure. Unlike geometric GNNs like GemNet (Gasteiger et al., 2021), 3rd-order features like angles are not necessary, due to our novel triplet interaction (as shown by (Li et al., 2024)). Additionally, unlike UniMol+ (Lu et al., 2023), no initial 3D structure is necessary for our approach, which is a significant advantage since the whole inference pipeline can be pure neural network-based and highly scalable. We predict binned distances instead of continuous values since distance distributions are often multimodal and skewed, and thus only fully captured by the model with a cross-entropy loss. This allows us to predict the most probable distance, i.e. the mode, which is more accurate than the mean or median predictions. Quantization noise due to binning does not affect the downstream task predictor which is robust to noise.
@@ 3.3 Stochastic Inference
| Val. MAE | Test-dev MAE | ||
| Model | #Param | (meV) | (meV) |
@@
| Val. Avg. | Test Avg. | |||
| Split | Uni-Mol+ | TGT-At | Uni-Mol+ | TGT-At |
| ID MAE | 379.5 | 381.3 | 374.5 | 379.6 |
| ID EwT | 11.15 | 11.15 | 11.29 | 11.50 |
| OOD Ads. MAE | 452.6 | 445.4 | 476.0 | 471.8 |
| OOD Ads. EwT | 6.71 | 6.87 | 6.05 | 5.70 |
| OOD Cat. MAE | 401.1 | 391.7 | 398.0 | 399.0 |
| OOD Cat. EwT | 9.90 | 10.47 | 9.53 | 9.84 |
| OOD Both MAE | 402.1 | 393.6 | 408.6 | 408.4 |
| OOD Both EwT | 6.68 | 6.80 | 6.06 | 6.17 |